An object falls 10.0m from rest before entering some water. Assuming negligible air resistance, what is the time taken to reach the water and the speed with which the object reaches the water? time/ ms speed/ ms–1 A 1.02 10.0 B 1.02 14.0 C 1.43 10.0 D 1.43 14.0


1
Expert's answer
2021-09-19T10:52:36-0400

"v_0 =0;g=9.8\frac{m}{s^2};s=10m"

"s = v_0t+\frac{gt^2}{2}"

"s = \frac{gt^2}{2}"

"t=\sqrt{\frac{2s}{g}}=\sqrt{\frac{2*10}{9.8}}\approx1.43s"

"v = v_0t+gt=9.8*1.43\approx 14 m"


"\text{Answer: }D\ 1.43 \ 14.0"